OpenStudy (anonymous):

Please help I am confused with this question How many grams of water can be produced when 11.7 moles of ethane (C2H6) react with excess oxygen gas? Unbalanced equation: C2H6 + O2 → CO2 + H2O Show, or explain, all of your work along with the final answer.

OpenStudy (anonymous):

I already balanced it too

OpenStudy (anonymous):

2C2H6 + 7O2 -> 4CO2 + 6H2O

OpenStudy (anonymous):

ps please reply as asap

OpenStudy (aaronq):

Use the stoichiometric coefficients to find moles produced. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) -------------------------------------------------------- To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles.

OpenStudy (anonymous):

could you help set up the equation with me

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

i can't. you should be able to after reading what i posted.

OpenStudy (anonymous):

well thanks lol

OpenStudy (aaronq):

no problem